OpenStudy (anonymous):

Simplify and write in terms of sine. (csc^2Ɵ - cot^2Ɵ) sec^2Ɵ

OpenStudy (anonymous):

Note that \(\csc^2\theta-\cot^2\theta = \dfrac{1}{\sin^2\theta}-\dfrac{\cos^2\theta}{\sin^2\theta} = \dfrac{1-\cos^2\theta}{\sin^2\theta}\), right? \(1-\cos^2\theta\) looks like it can be simplified... maybe it has something to do with \(\sin^2\theta+\cos^2\theta=1\)?

OpenStudy (anonymous):

\[1-\cos^2\theta = \sin^2\theta \] @micahwood50

OpenStudy (anonymous):

Right, so you can replace that to \(\sin^2\theta\), then simplify ahead

OpenStudy (anonymous):

You should see that \(\csc^2\theta-\cot^2\theta\) is equal to one.

OpenStudy (anonymous):

ohhh I get it so it will be \[\frac{ \sin^2\theta }{ \sin^2\theta }\] @micahwood50

OpenStudy (anonymous):

Right, hence 1

OpenStudy (anonymous):

So you have this: \((\csc^2\theta-\cot^2\theta)\sec^2\theta ~\Rightarrow~(1)\sec^2\theta~\Rightarrow~?\)

OpenStudy (anonymous):

So now I have to write \[\sec^2\theta \] in terms of sine? @micahwood50

OpenStudy (anonymous):

Isn't \[\sec^2 \theta = \frac{ 1 }{ \cos^2 \theta }\] @micahwood50

OpenStudy (anonymous):

oh right, i forgot about writing in term of sine. just rewrite \(\dfrac{1}{\cos^2\theta}\). Again, \(\sin^2\theta+cos^2\theta=1\)